# 1]Fundamental AI Research (FAIR) at Meta

### v1

@@ Appendix C Inference efficiency
Figure C.11: Predicted phonon band structure and density of states (DOS) of Si (diamond structure), CsCl (CsCl structure), AlN (wurtzite structure) using eqV2 S DeNS (direct-force prediction) at different displacement values. DFT baseline is taken from the PBE MDR dataset Loew et al. (2024) calculated using a displacement of 0.01 Å
Figure C.12: Predicted phonon band structure and density of states (DOS) of Si (diamond structure), CsCl (CsCl structure), AlN (wurtzite structure) using eqV2 S DeNS (direct-force prediction with a zero net force constraint) at different displacement values. DFT baseline is taken from the PBE MDR dataset Loew et al. (2024) calculated using a displacement of 0.01 Å
@@ Appendix D Hyper-parameters
Hyper-parameters used for model training are shown in Table 5 We train all models using a per-atom energy MAE loss, a force $l_{2}$ loss, and a stress MAE loss. For direct-force models or direct-force pre-training, we use the same decomposed loss as described in  Barroso-Luque et al. 2024. The eSEN-30M model trained on MPTrj uses Denoising Non-equilibrium Structures (DeNS) (Liao et al., 2024) with a noising probablity of $0.5$, a standard deviation of $0.1$ for the added Gaussian noise, and $10$ for the DeNS loss coefficient during direct-force pre-training. DeNS is not used during conservative fine-tuning. In our ablation study for maximum neighbor limit we used $30$ as the limit.
Table 5: Hyper-parameters for eSEN trained on MPTrj/SPICE-MACE-OFF. *eSEN-30M on MPTrj was trained for 60 epochs using direct-force pre-training and 40 epochs of conserved fine-tuning.
Hyper-parameters SPICE-3.2M SPICE-6.5M MPTrj-3.2M MPTrj-6.5M MPTrj-30M Number of eSEN layer blocks $2$ $4$ $2$ $4$ $10$ Maximum degree $L_{\mathrm{max}}$ $2$ $2$ $2$ $2$ $3$ Maximum order $M_{\mathrm{max}}$ $2$ $2$ $2$ $2$ $2$ Number of channels $N_{\mathrm{channel}}$ $128$ $128$ $128$ $128$ $128$ Radial basis function Bessel Bessel Gaussian Gaussian Gaussian Number of radial basis functions $10$ $10$ $10$ $10$ $10$ Cutoff radius (Å) $5$ $5$ $6$ $6$ $6$ Batch size 128 128 512 512 512 Optimizer AdamW AdamW AdamW AdamW AdamW Learning rate scheduling Cosine Cosine Cosine Cosine Cosine Warmup epochs $0.1$ $0.01$ $0.1$ $0.1$ $0.1$ Warmup factor $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ Maximum learning rate $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ Number of epochs $100$ $100$ $100$ $100$ $60+40$* Gradient clipping norm threshold $100$ $100$ $100$ $100$ $100$ Model EMA decay $0.999$ $0.999$ $0.999$ $0.999$ $0.999$ Weight decay $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ Energy loss coefficient $10$ $10$ $1$ $1$ $20$ Force loss coefficient $20$ $20$ $10$ $10$ $20$ Stress loss coefficient - - $100$ $100$ $5$

### v2

@@ Appendix C Inference efficiency
Figure C.11: Predicted phonon band structure and density of states (DOS) of Si (diamond structure), CsCl (CsCl structure), AlN (wurtzite structure) using eqV2-S-DeNS (direct-force prediction) at different displacement values. DFT baseline is taken from the PBE MDR dataset Loew et al. (2024) calculated using a displacement of 0.01 Å
Figure C.12: Predicted phonon band structure and density of states (DOS) of Si (diamond structure), CsCl (CsCl structure), AlN (wurtzite structure) using eqV2-S-DeNS (direct-force prediction with a zero net force constraint) at different displacement values. DFT baseline is taken from the PBE MDR dataset Loew et al. (2024) calculated using a displacement of 0.01 Å
@@ Appendix D Hyper-parameters
Hyper-parameters used for model training are shown in Table 6. We train all models using a per-atom energy MAE loss, a force $l_{2}$ loss, and a stress MAE loss. For direct-force models or direct-force pre-training, we use the same decomposed loss as described in  Barroso-Luque et al. 2024. The eSEN-30M model trained on MPTrj uses Denoising Non-equilibrium Structures (DeNS) (Liao et al., 2024) with a noising probablity of $0.5$, a standard deviation of $0.1$ for the added Gaussian noise, and $10$ for the DeNS loss coefficient during direct-force pre-training. DeNS is not used during OMat24 training or conservative fine-tuning. In our ablation study for maximum neighbor limit we used $30$ as the limit.
Table 6: Hyper-parameters for eSEN variants reported in this paper. *eSEN-30M on MPTrj was trained for 60 epochs using direct-force pre-training and 40 epochs of conserved fine-tuning. eSEN-30M-OMat was trained for 2 epochs using direct-force pre-training and 2 epochs of conserved fine-tuning. The eSEN-30M-OAM model starts from the eSEN-30M-OMat model, and was finetuned for 1 epoch on a dataset constructed by combining the sAlex training dataset and 8 copies of the MPTrj training dataset.
Hyper-parameters SPICE-3.2M SPICE-6.5M MPTrj-3.2M MPTrj-6.5M MPTrj-30M OMat-30M OAM Fine-tuning Number of eSEN layer blocks $2$ $4$ $2$ $4$ $10$ $10$ $10$ Maximum degree $L_{\mathrm{max}}$ $2$ $2$ $2$ $2$ $3$ $3$ $3$ Maximum order $M_{\mathrm{max}}$ $2$ $2$ $2$ $2$ $2$ $2$ $2$ Number of channels $N_{\mathrm{channel}}$ $128$ $128$ $128$ $128$ $128$ $128$ $128$ Radial basis function Bessel Bessel Gaussian Gaussian Gaussian Gaussian Gaussian Number of radial basis functions $10$ $10$ $10$ $10$ $10$ $64$ $64$ Cutoff radius (Å) $5$ $5$ $6$ $6$ $6$ $6$ $6$ Batch size 128 128 512 512 512 512 256 Optimizer AdamW AdamW AdamW AdamW AdamW AdamW AdamW Learning rate scheduling Cosine Cosine Cosine Cosine Cosine Cosine Cosine Warmup epochs $0.1$ $0.01$ $0.1$ $0.1$ $0.1$ $0.1$ $0.1$ Warmup factor $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ $0.2$ Maximum learning rate $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $4\times 10^{-4}$ $2\times 10^{-4}$ Number of epochs $100$ $100$ $100$ $100$ $60+40$* $2+2$* $1$ Gradient clipping norm $100$ $100$ $100$ $100$ $100$ $100$ $100$ Model EMA decay $0.999$ $0.999$ $0.999$ $0.999$ $0.999$ $0.999$ $0.999$ Weight decay $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ $1\times 10^{-3}$ Energy loss coefficient $10$ $10$ $1$ $1$ $20$ $20$ $20$ Force loss coefficient $20$ $20$ $10$ $10$ $20$ $20$ $20$ Stress loss coefficient - - $100$ $100$ $5$ $5$ $5$